### v1

@@ 6 Conclusion
This paper considers a practically motivated variant of the canonical problem of online fair allocation wherein a decision-maker has a budget of perishable resources to allocate fairly and efficiently over a fixed time horizon. Our main insight is that perishability fundamentally impacts the envy-efficiency trade-off derived for the no-perishing setting: while a decision-maker can arbitrarily sacrifice on envy in favor of efficiency in this latter setting, this is no longer the case when there is uncertainty around items’ perishing times. We derive strong lower bounds to formalize this insight, which are a function of both the quality of the decision-maker’s prediction over perishing times, as well as the inherent aggressiveness of the perishing process. We moreover design an algorithm that achieves these lower bounds; this algorithm relies on the construction of a baseline allocation that accounts for the unavoidable spoilage incurred by any online algorithm. From a technical perspective, the main challenge that the perishing setting presents is that the uncertainty around the quantity of resources that spoil in the future is endogenous, in contrast to the exogenous uncertainty on the number of arrivals in the classical setting. Deriving tight bounds on spoilage (both for our lower bounds as well as in the design of our algorithm) relied on the “slow allocation” construction, which rendered the highly coupled process amenable to tractable analysis. Finally, our numerical experiments demonstrate our algorithm’s strong performance against state-of-the-art perishing-agnostic benchmarks.
@@
In terms of future directions, our work identified offset-expiry as a necessary condition for which the classical notion of envy-freeness is even meaningful. While our algorithm performs well numerically in “low-probability” offset-expiring settings, relaxing this assumption in theory remains an interesting open question. Though we conjecture that the slow allocation construction will remain a useful tool in more aggressive perishing settings, the philosophical question of how to define more appropriate notions of envy is likely the more important one. In addition to this, our model assumes that the decision-maker allocates items according to a fixed allocation schedule. Though our results do not require that the perishing distribution be memoryless, allowing for time-varying / adaptive allocation schedules, though less practical, would improve our algorithm’s performance in non-memoryless settings. This relates back to the question of deriving theoretical insights into the structure of optimal allocation schedules. Finally, though this paper considered exogenous depletion of the budget, a natural practical extension is one wherein $B$ evolves stochastically, accounting for external donations independent of the allocations made by the algorithm.
Acknowledgments

### v2

@@ 6 Conclusion
This paper considers a practically motivated variant of the canonical problem of online fair allocation wherein a decision-maker has a budget of perishable resources to allocate fairly and efficiently over a fixed time horizon. Our main insight is that perishability fundamentally impacts the envy-efficiency trade-off derived for the no-perishing setting: while a decision-maker can arbitrarily sacrifice on envy in favor of efficiency in this latter setting, this is no longer the case when there is uncertainty around items’ perishing times. We derive strong lower bounds to formalize this insight, which are a function of both the quality of the decision-maker’s prediction over perishing times, as well as the inherent aggressiveness of the perishing process. We moreover design an algorithm that achieves these lower bounds; this algorithm relies on the construction of a baseline allocation that accounts for the unavoidable spoilage incurred by any online algorithm. From a technical perspective, the main challenge that the perishing setting presents is that the uncertainty around the quantity of resources that spoil in the future is endogenous, in contrast to the exogenous uncertainty on the number of arrivals in the classical setting. Deriving tight bounds on spoilage (both for our lower bounds as well as in the design of our algorithm) relied on the “slow allocation” construction, which rendered the highly coupled process amenable to tractable analysis. Our numerical experiments also demonstrate our algorithm’s strong performance against state-of-the-art perishing-agnostic benchmarks.
From a practical perspective, our results provide concrete, intuitive guidance as to the composition of perishable donations that a food bank should accept in order to achieve a fair allocation (in the proportional sense). At a high level, the optimal mix of goods should ensure that cumulative perishing always lags behind cumulative demand. Moreover, we uncover the important — and counterintuitive — insight that guaranteeing a fair allocation in the presence of perishable goods may require the decision-maker to allocate significantly less than the static proportional benchmark. Finally, our numerical results suggest that an allocation schedule that ranks items according to high-probability lower bounds on their perishing times is a robust choice that appropriately hedges against the inherent variability of the perishing process. In particular, we observe that this allocation schedule outperforms natural alternatives such as ordering items according to expected perishing time.
@@
In terms of future directions, our model assumes that the decision-maker allocates items according to a fixed allocation schedule. Though our results do not require that the perishing distribution be memoryless, allowing for time-varying / adaptive allocation schedules, though less practical, would improve our algorithm’s performance in non-memoryless settings. This relates back to the question of deriving theoretical insights into the structure of optimal allocation schedules. Finally, though this paper considered exogenous depletion of the budget, a natural practical extension is one wherein $B$ evolves stochastically, accounting for external donations independent of the allocations made by the algorithm.
[…]